OpenStudy (anonymous):

I=PRT; I=4500, P=25,000,R=0.06

OpenStudy (anonymous):

4500=(25000)(0.06)(T)

OpenStudy (anonymous):

HEY U HAVE TO SOLVE IT DAT NOT GOOD ANSWER BRO

OpenStudy (anonymous):

lol! ok T= 4500/(25000*0.06) T=4500/1500 T=5

OpenStudy (anonymous):

U MISS THE ANSWER THE ANSWER IS 3

OpenStudy (anonymous):

yes it is 3 sry